Write the correct Lewis structure of product and show the step-by-step solution on how you came up with the structure.

SOCl2- primarily used as a chlorinating reagent.


1
Expert's answer
2021-12-17T06:28:02-0500

Lewis structure of SOCl"_2"






Number of valence electrons

"=6+6+7(2)"

"=26"


Adding dots




Formal charge"=" "\begin{matrix}
 Valence \>electrons- Non\>Bonding \>val\>electrons&-\frac{Bonding\>electrons}{2}
 
\end{matrix}"


"Cl=7-6-\frac{2}{2}=0\\S=6-2-\frac{6}{2}=+1\\O=6-6-\frac{2}{2}=-1"



Improved structure





Formal charges;

"Cl=7-6-\frac{2}{2}=0\\S=6-2-\frac{8}{2}=0\\O=6-4-\frac{4}{2}=0"


This is the best structure